# A uniform kernel trick for high and infinite-dimensional two-sample problems

### v2

@@ Abstract
We use a suitable version of the so-called “kernel trick” to devise two-sample tests, especially focussed on high-dimensional and functional data. Our proposal entails a simplification of the practical problem of selecting an appropriate kernel function. Specifically, we apply a uniform variant of the kernel trick which involves the supremum within a class of kernel-based distances. We obtain the asymptotic distribution of the test statistic under the null and alternative hypotheses. The proofs rely on empirical processes theory, combined with the delta method and Hadamard directional differentiability techniques, and functional Karhunen-Loève-type expansions of the underlying processes. This methodology has some advantages over other standard approaches in the literature. We also give some experimental insight into the performance of our proposal compared to other kernel-based approaches (the original proposal by Borgwardt et al. (2006) and some variants based on splitting methods) as well as tests based on energy distances Rizzo and Székely (2017).
@@ 1 Introduction: an overview
We focus on statistical problems where, essentially, the aim is to properly separate data coming from two different populations; this is the case of binary supervised classification and two-sample testing problems. In such situations, the kernel trick is a common paradigm. In a few words, the standard multivariate version (i.e., with data in $\mathbb{R}^{d}$) of the kernel trick lies in separating the data in both populations using a symmetric non-negative definite “kernel function”. The values of the kernel can be seen as the inner product of transformed versions of the original observations in a different (usually higher-dimensional) space. It is expected that the groups can be better distinguished in the new final space; see Scholköpf and Smola (2018). We are particularly interested in those situations in which the available data are high-dimensional or even functional (thus, infinite-dimensional). In such cases, the strategy of mapping the data into a higher-dimensional space does not seem to be so compelling. Still, the kernel trick remains meaningful in a sort of “second generation” version, whose point is to take the data to a more comfortable and flexible space. In this new space, the statistical methodology might be mathematically more tractable, and more easily implemented and interpreted. To be more precise, a probability distribution $\operatorname{P}$ on the sample space $\mathcal{X}$ is replaced with a function
@@
in an appropriate space of “nice functions” defined by means of the kernel $k$. In this way, the distance between two probability measures is computed in terms of the metric in the functional space. As a matter of fact, one of the most appealing proposals in this direction relies on kernel-based distances, expressed in terms of the embedding transformation $\mu_{\operatorname{P}}$ in (1); see Borgwardt et al. (2006).
@@
Further, a parameter-dependent method might be an obstacle for practitioners who are often reluctant to use procedures depending on auxiliary, hard-to-interpret parameters. We thus find here a particular instance of the trade-off between power and applicability: as stated in Tukey (1959), the practical power of a statistical procedure is defined as “the product of the mathematical power by the probability that the procedure will be used” (Tukey credits to Churchill Eisenhart for this idea). From this perspective, our proposal can be viewed as an attempt to make kernel-based homogeneity tests more usable by getting rid of the tuning parameter(s). Roughly speaking, the idea that we propose to avoid selecting a specific value of $\lambda$ within the family $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ is to take the supremum over the set of parameters $\Lambda$ of the resulting family of kernel-distances. We call this approach the uniform kernel trick, as we map the data into many functional spaces at the same time and use, as test statistic, the supremum of the corresponding kernel distances. We believe that this methodology could be successfully applied as well in supervised classification, though this topic is not considered in this work.
@@
The energy distance (see the review by Rizzo and Székely (2017)) and the associated distance covariance, as well as kernel distance, represent a step forward in this direction since they can be calculated with relative ease for high-dimensional distributions. In Fukumizu et al. (2013) the relationships among these metrics in the context of hypothesis testing are discussed. In this paper we consider an extension, as well as an alternative mathematical approach, for the two-sample test in Borgwardt et al. (2006). These authors show that kernel-based procedures perform better than other more classical approaches when dimension grows, although they are strongly dependent on the choice of the kernel parameter.
@@
To present the contributions of this paper, we briefly refer to some important, mutually related, technical notions. As emphasized in Berlinet and Thomas-Agnan (2011), Reproducing Kernel Hilbert Spaces (RKHS in short) provide an excellent environment to construct helpful transformations in several statistical problems. Given a topological space $\mathcal{X}$ (in many applications $\mathcal{X}$ is a subset of a Hilbert space), a kernel $k$ is a real non-negative semidefinite symmetric function on $\mathcal{X}\times\mathcal{X}$. The RKHS associated with $k$, denoted in the following by $\mathcal{H}_{k}$, is the Hilbert space generated by finite linear combinations of type $\sum_{j}\alpha_{j}\,k\left(x_{j},\cdot\right)$; see Section 2 for additional details. Let $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$ be the set of (Borel) probability measures on $\mathcal{X}$. Under mild assumptions on $k$, the functions in $\mathcal{H}_{k}$ are measurable and $\operatorname{P}$-integrable, for each $\operatorname{P}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$. Moreover, it can be checked that the function $\mu_{\operatorname{P}}$ in (1) belongs to $\mathcal{H}_{k}$. The transformation $\operatorname{P}\mapsto\mu_{\operatorname{P}}$ from $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$ to $\mathcal{H}_{k}$ is called the (kernel) mean embedding; see Fukumizu et al. (2013) and (Berlinet and Thomas-Agnan, 2011, Chapter 4). The mean embedding of $\operatorname{P}$ can be viewed as a smoothed version of the distribution of $\operatorname{P}$ through the kernel $k$ within the RKHS. This is evident when $\operatorname{P}$ is absolutely continuous with density $f$ and $k(x,y)=K(x-y)$, for some real function $K$. In this situation, $\mu_{\operatorname{P}}$ is the convolution of $f$ and $K$. On the other hand, mean embeddings appear, under the name of potential functions, in some other mathematical fields (such as functional analysis); see (El-Fallah et al., 2014, p. 15). The kernel distance between $\operatorname{P}$ and $\operatorname{Q}$ in $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$ is
@@
where $\|\cdot\|_{\mathcal{H}_{k}}$ stands for the norm in $\mathcal{H}_{k}$ and $(\operatorname{P}-\operatorname{Q})\otimes(\operatorname{P}-\operatorname{Q})$ denotes the product (signed) measure on $\mathcal{X}^{2}=\mathcal{X}\times\mathcal{X}$. Therefore, $d_{k}(\operatorname{P},\operatorname{Q})$ is the RKHS distance between the mean embeddings of the corresponding probability measures. Kernel distances were popularized in machine learning as tools to tackle several relevant statistical problems, such as homogeneity tests Borgwardt et al. (2006), independence Fukumizu et al. (2007), test of conditional independence Fukumizu et al. (2008) and density estimation Fukumizu et al. (2011). The key idea behind this methodology can be seen as a particular case of the fruitful kernel trick paradigm.
@@
We consider a family of kernels $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$, where $\Lambda$ is certain parametric space. For the Gaussian kernel in (2), $\Lambda=(0,\infty)$, but in general $\lambda$ could be a multidimensional parameter, as in the case of Matérn kernels or inverse quadratic kernels; see (Sriperumbudur, 2016, p. 1846). Each $k_{\lambda}$ has an associated RKHS, $\mathcal{H}_{k,\lambda}$ (endowed with its intrinsic norm $\|\cdot\|_{\mathcal{H}_{k,\lambda}}$), and the corresponding probability distance $d_{k,\lambda}$. For $\operatorname{P},\operatorname{Q}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$, we want to test $\operatorname{H}_{0}:\operatorname{P}=\operatorname{Q}$ using the distances within the collection $\left\{d_{k,\lambda}:\lambda\in\Lambda\right\}$. The current theoretical framework does not support the automatic (data-driven) choice of $\lambda\in\Lambda$, since the asymptotic theory is mainly developed for a fixed kernel, corresponding to a specific value of $\lambda$. However, the choice of $\lambda$ is a non-trivial and sensitive issue with no obvious best solution, and which might affect the test performance.
There are various interesting proposals to deal with this problem in practice: the median heuristic of Borgwardt et al. (2006); sample-splitting and optimization methods in Balakrishnan et al. (2012) and Gretton et al. (2020); and aggregation methods such as Balakrishnan et al. (2012). In this paper we explore an alternative to avoid making a parametric decision or splitting the data set. Our proposal can be included within the aggregative methods: we combine the information provided by different kernels by taking the supremum over the induced kernel metrics. Specifically, we use the metric that “best separates” $\operatorname{P}$ and $\operatorname{Q}$, that is, the supremum of all kernel distances given by
@@
The supremum kernel distance (4) entails several advantages and some mathematical challenges: First, the kernel selection problem is considerably simplified and solved in a natural way. Additionally, the approach is general enough to be applied in infinite-dimensional settings as FDA. This is interesting since in FDA there are only a few homogeneity tests in the literature. Some of them have been developed in the setting of ANOVA models (involving several samples) under homoscedasticity (equal covariance operators of the involved processes) and Gaussian assumptions. Hence, the current methodologies amount to testing the null hypothesis of equal means in all the populations; see, e.g., Cuevas et al. (2004) for an early contribution and Zhang (2014) for a broader perspective. Our proposal is therefore quite related to more general approaches, not requiring any homoscedasticity assumption and still valid for a FDA framework. Examples of such similar tests are Ghosh et al. (2016) and Hall and Van Keilegom (2007), as well as the random projections-based methodology in Cuesta-Albertos et al. (2007).
The inclusion of the supremum in (4) represents an additional difficulty. The asymptotic properties of the test statistic based on (4) are derived by following a different strategy from that of Borgwardt et al. (2006) and later works. The methodology proposed here allows us to cope with the supremum and applies directly to the case of unequal sample sizes. In short, our approach can be described as follows: First, we consider plug-in estimators of the kernel distances, obtained by replacing the unknown distributions by their empirical counterparts. Then, we use the powerful theory of empirical processes together with some recent results on the differentiability of the supremum (see Cárcamo et al. (2020)) and functional Karhunen-Loève expansions of the underlying processes. These developments entail several technical difficulties from the mathematical point of view. However, they are worthwhile since they allow us to analyze the asymptotic behavior, under both the null and the alternative hypothesis, of the two-sample test based on (4).
@@ 2 Preliminaries
The theory of RKHS plays a relevant role in this paper. This is a classical and well-known topic; see (Janson, 1997, Appendix F) for a brief account of the RKHS theory and Berlinet and Thomas-Agnan (2011) or Eubank and Hsing (2015) for a statistical perspective. Hence, we only mention what is strictly necessary for later use. Let $\mathcal{X}$ be a topological space and $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ a kernel, that is, a symmetric and positive semi-definite function. Let us consider $\mathcal{H}^{0}_{k}$, the pre-Hilbert space of all finite linear combinations $g(\cdot)=\sum_{i=1}^{n}\alpha_{i}\,k\left(x_{i},\cdot\right)$ (with $\alpha_{i}\in\mathbb{R}$, $n\in\mathbb{N}$ and $x_{i}\in\mathcal{X}$), endowed with the inner product
@@
The RKHS $\mathcal{H}_{k}$ is defined as the completion of $\mathcal{H}_{k}^{0}$; see (Berlinet and Thomas-Agnan, 2011, Chapter 1). The inner product $\langle\cdot,\cdot\rangle_{k}$ in $\mathcal{H}_{k}$ is obtained through (5) in such a way that bilinearity is preserved. A key property of RKHS is the so-called reproducing property:
@@
Sufficient conditions guaranteeing the injectivity of the mean embedding transformation can be found in Fukumizu et al. (2011). Note that in (7) (and what follows) we use the standard notation in empirical processes theory: $\operatorname{P}(f)$ (or simply $\operatorname{P}f$) stands for the mathematical expectation of $f$ with respect to $\operatorname{P}$. The existence of the mean embedding implies that the kernel distance in (3), as well as the supremum kernel distance in (4), are well-defined. Indeed, they are integral probability metrics; see Müller (1997). To see this, let us consider the unit ball of $\mathcal{H}_{k}$, that is,
@@
where $(a)$ follows from the definition of mean embedding (1), $(b)$ from Pettis integrability, and $(c)$ from the reproducing property (6); see also (Borgwardt et al., 2012, Lemma 4). Thus, the kernel distance (3) is the integral probability metric generated by the class $\mathcal{F}_{k}$ in (9). Therefore, the supremum kernel distance (4) admits the alternative representation
@@
From the characterizations as integral probability metrics in (10) and (11), we conclude that $d_{k}$ and $d_{k,\Lambda}$ satisfy the properties of a pseudo-metric (non-negativeness, symmetry, triangular property). However, to ensure the identifiability property of a metric $d$ (i.e., $d(\operatorname{P},\operatorname{Q})=0$ if and only if $\operatorname{P}=\operatorname{Q}$) additional conditions are needed. It can be checked that when $\mathcal{X}=\mathbb{R}^{d}$, identifiability is satisfied for the usual kernels (such as the Gaussian kernel in (2)). However, when $\mathcal{X}$ is infinite-dimensional this type of results are more complicated; see Duncan and Wynne (2022) for a deep study of this topic for the Gaussian kernel (2). More details can also be found in Fukumizu et al. (2010) and Fukumizu et al. (2011).
@@
As a supremum over a class of functions is involved in (12), the theory of empirical processes comes into play naturally. Given a collection of functions $\mathcal{F}$, we recall that the $\mathcal{F}$-indexed empirical process (associated to $\operatorname{P}$) is $\mathbb{G}_{n}^{\operatorname{P}}=\sqrt{n}\,\left(\mathbb{P}_{n}-\operatorname{P}\right)$. The class $\mathcal{F}$ is called P-Donsker if $\mathbb{G}_{n}^{\operatorname{P}}\rightsquigarrow\mathbb{G}_{\operatorname{P}}$ in $\ell^{\infty}(\mathcal{F})$, the space of bounded real functionals defined on $\mathcal{F}$ with the supremum norm; see van der Vaart and Wellner (2023). Here, $\rightsquigarrow$ stands for weak convergence in $\ell^{\infty}(\mathcal{F})$ and $\mathbb{G}_{\operatorname{P}}$ is a $\operatorname{P}$-Brownian bridge, that is, a zero-mean Gaussian process with covariance function

### v3

@@ Abstract
We use a suitable version of the so-called “kernel trick” to devise two-sample tests, especially focussed on high-dimensional and functional data. Our proposal entails a simplification of the practical problem of selecting an appropriate kernel function. Specifically, we apply a uniform variant of the kernel trick which involves the supremum within a class of kernel-based distances. We obtain the asymptotic distribution of the test statistic under the null and alternative hypotheses. The proofs rely on empirical processes theory, combined with the delta method and Hadamard directional differentiability techniques, and functional Karhunen-Loève-type expansions of the underlying processes. This methodology has some advantages over other standard approaches in the literature. We also give some experimental insight into the performance of our proposal compared to other kernel-based approaches (the original proposal by [10] and some variants based on splitting methods) as well as tests based on energy distances [41].
@@ 1 Introduction: an overview
We focus on statistical problems where, essentially, the aim is to properly separate data coming from two different populations; this is the case of binary supervised classification and two-sample testing problems. In such situations, the kernel trick is a common paradigm. In a few words, the standard multivariate version (i.e., with data in $\mathbb{R}^{d}$) of the kernel trick lies in separating the data in both populations using a symmetric non-negative definite “kernel function”. The values of the kernel can be seen as the inner product of transformed versions of the original observations in a different (usually higher-dimensional) space. It is expected that the groups can be better distinguished in the new final space; see [43].
We are particularly interested in those situations in which the available data are high-dimensional or even functional (thus, infinite-dimensional). In such cases, the strategy of mapping the data into a higher-dimensional space does not seem to be so compelling. Still, the kernel trick remains meaningful in a sort of “second generation” version, whose point is to take the data to a more comfortable and flexible space. In this new space, the statistical methodology might be mathematically more tractable, and more easily implemented and interpreted. To be more precise, a probability distribution $\operatorname{P}$ on the sample space $\mathcal{X}$ is replaced with a function
@@
in an appropriate space of “nice functions” defined by means of the kernel $k$. In this way, the distance between two probability measures is computed in terms of the metric in the functional space. As a matter of fact, one of the most appealing proposals in this direction relies on kernel-based distances, expressed in terms of the embedding transformation $\mu_{\operatorname{P}}$ in (1); see [10].
@@
Further, a parameter-dependent method might be an obstacle for practitioners who are often reluctant to use procedures depending on auxiliary, hard-to-interpret parameters. We thus find here a particular instance of the trade-off between power and applicability: as stated in [47], the practical power of a statistical procedure is defined as “the product of the mathematical power by the probability that the procedure will be used” (Tukey credits to Churchill Eisenhart for this idea). From this perspective, our proposal can be viewed as an attempt to make kernel-based homogeneity tests more usable by getting rid of the tuning parameter(s). Roughly speaking, the idea that we propose to avoid selecting a specific value of $\lambda$ within the family $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$ is to take the supremum over the set of parameters $\Lambda$ of the resulting family of kernel-distances. We call this approach the uniform kernel trick, as we map the data into many functional spaces at the same time and use, as test statistic, the supremum of the corresponding kernel distances. We believe that this methodology could be successfully applied as well in supervised classification, though this topic is not considered in this work.
@@
The energy distance (see the review by [41]) and the associated distance covariance, as well as kernel distance, represent a step forward in this direction since they can be calculated with relative ease for high-dimensional distributions. In [28] the relationships among these metrics in the context of hypothesis testing are discussed. In this paper we consider an extension, as well as an alternative mathematical approach, for the two-sample test in [10]. These authors show that kernel-based procedures perform better than other more classical approaches when dimension grows, although they are strongly dependent on the choice of the kernel parameter.
@@
To present the contributions of this paper, we briefly refer to some important, mutually related, technical notions. As emphasized in [6], Reproducing Kernel Hilbert Spaces (RKHS in short) provide an excellent environment to construct helpful transformations in several statistical problems. Given a topological space $\mathcal{X}$ (in many applications $\mathcal{X}$ is a subset of a Hilbert space), a kernel $k$ is a real non-negative semidefinite symmetric function on $\mathcal{X}\times\mathcal{X}$. The RKHS associated with $k$, denoted in the following by $\mathcal{H}_{k}$, is the Hilbert space generated by finite linear combinations of type $\sum_{j}\alpha_{j}\,k\left(x_{j},\cdot\right)$; see Section 2 for additional details.
Let $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$ be the set of (Borel) probability measures on $\mathcal{X}$. Under mild assumptions on $k$, the functions in $\mathcal{H}_{k}$ are measurable and $\operatorname{P}$-integrable, for each $\operatorname{P}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$. Moreover, it can be checked that the function $\mu_{\operatorname{P}}$ in (1) belongs to $\mathcal{H}_{k}$. The transformation $\operatorname{P}\mapsto\mu_{\operatorname{P}}$ from $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$ to $\mathcal{H}_{k}$ is called the (kernel) mean embedding; see [28] and [6, Chapter 4]. The mean embedding of $\operatorname{P}$ can be viewed as a smoothed version of the distribution of $\operatorname{P}$ through the kernel $k$ within the RKHS. This is evident when $\operatorname{P}$ is absolutely continuous with density $f$ and $k(x,y)=K(x-y)$, for some real function $K$. In this situation, $\mu_{\operatorname{P}}$ is the convolution of $f$ and $K$. On the other hand, mean embeddings appear, under the name of potential functions, in some other mathematical fields (such as functional analysis); see [20, p. 15].
The kernel distance between $\operatorname{P}$ and $\operatorname{Q}$ in $\mathcal{M}_{\operatorname{p}}(\mathcal{X})$ is
@@
where $\|\cdot\|_{\mathcal{H}_{k}}$ stands for the norm in $\mathcal{H}_{k}$ and $(\operatorname{P}-\operatorname{Q})\otimes(\operatorname{P}-\operatorname{Q})$ denotes the product (signed) measure on $\mathcal{X}^{2}=\mathcal{X}\times\mathcal{X}$. Therefore, $d_{k}(\operatorname{P},\operatorname{Q})$ is the RKHS distance between the mean embeddings of the corresponding probability measures. Kernel distances were popularized in machine learning as tools to tackle several relevant statistical problems, such as homogeneity tests [10], independence [23], test of conditional independence [24] and density estimation [27]. The key idea behind this methodology can be seen as a particular case of the fruitful kernel trick paradigm.
@@
We consider a family of kernels $\left\{k_{\lambda}:\lambda\in\Lambda\right\}$, where $\Lambda$ is certain parametric space. For the Gaussian kernel in (2), $\Lambda=(0,\infty)$, but in general $\lambda$ could be a multidimensional parameter, as in the case of Matérn kernels or inverse quadratic kernels; see [46, p. 1846]. Each $k_{\lambda}$ has an associated RKHS, $\mathcal{H}_{k,\lambda}$ (endowed with its intrinsic norm $\|\cdot\|_{\mathcal{H}_{k,\lambda}}$), and the corresponding probability distance $d_{k,\lambda}$. For $\operatorname{P},\operatorname{Q}\in\mathcal{M}_{\operatorname{p}}(\mathcal{X})$, we want to test $\operatorname{H}_{0}:\operatorname{P}=\operatorname{Q}$ using the distances within the collection $\left\{d_{k,\lambda}:\lambda\in\Lambda\right\}$. The current theoretical framework does not support the automatic (data-driven) choice of $\lambda\in\Lambda$, since the asymptotic theory is mainly developed for a fixed kernel, corresponding to a specific value of $\lambda$. However, the choice of $\lambda$ is a non-trivial and sensitive issue with no obvious best solution, and which might affect the test performance.
There are various interesting proposals to deal with this problem in practice: the median heuristic of [10]; sample-splitting and optimization methods in [3] and [32]; and aggregation methods such as [3]. In this paper we explore an alternative to avoid making a parametric decision or splitting the data set. Our proposal can be included within the aggregative methods: we combine the information provided by different kernels by taking the supremum over the induced kernel metrics. Specifically, we use the metric that “best separates” $\operatorname{P}$ and $\operatorname{Q}$, that is, the supremum of all kernel distances given by
@@
The supremum kernel distance (4) entails several advantages and some mathematical challenges: First, the kernel selection problem is considerably simplified and solved in a natural way. Additionally, the approach is general enough to be applied in infinite-dimensional settings as FDA. This is interesting since in FDA there are only a few homogeneity tests in the literature. Some of them have been developed in the setting of ANOVA models (involving several samples) under homoscedasticity (equal covariance operators of the involved processes) and Gaussian assumptions. Hence, the current methodologies amount to testing the null hypothesis of equal means in all the populations; see, e.g., [16] for an early contribution and [52] for a broader perspective. Our proposal is therefore quite related to more general approaches, not requiring any homoscedasticity assumption and still valid for a FDA framework. Examples of such similar tests are [29] and [34], as well as the random projections-based methodology in [15].
The inclusion of the supremum in (4) represents an additional difficulty. The asymptotic properties of the test statistic based on (4) are derived by following a different strategy from that of [10] and later works. The methodology proposed here allows us to cope with the supremum and applies directly to the case of unequal sample sizes. In short, our approach can be described as follows: First, we consider plug-in estimators of the kernel distances, obtained by replacing the unknown distributions by their empirical counterparts. Then, we use the powerful theory of empirical processes together with some recent results on the differentiability of the supremum (see [14]) and functional Karhunen-Loève expansions of the underlying processes. These developments entail several technical difficulties from the mathematical point of view. However, they are worthwhile since they allow us to analyze the asymptotic behavior, under both the null and the alternative hypothesis, of the two-sample test based on (4).
@@ 2 Preliminaries
The theory of RKHS plays a relevant role in this paper. This is a classical and well-known topic; see [35, Appendix F] for a brief account of the RKHS theory and [6] or [21] for a statistical perspective. Hence, we only mention what is strictly necessary for later use. Let $\mathcal{X}$ be a topological space and $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ a kernel, that is, a symmetric and positive semi-definite function. Let us consider $\mathcal{H}^{0}_{k}$, the pre-Hilbert space of all finite linear combinations $g(\cdot)=\sum_{i=1}^{n}\alpha_{i}\,k\left(x_{i},\cdot\right)$ (with $\alpha_{i}\in\mathbb{R}$, $n\in\mathbb{N}$ and $x_{i}\in\mathcal{X}$), endowed with the inner product
@@
The RKHS $\mathcal{H}_{k}$ is defined as the completion of $\mathcal{H}_{k}^{0}$; see [6, Chapter 1]. The inner product $\langle\cdot,\cdot\rangle_{k}$ in $\mathcal{H}_{k}$ is obtained through (5) in such a way that bilinearity is preserved. A key property of RKHS is the so-called reproducing property:
@@
Sufficient conditions guaranteeing the injectivity of the mean embedding transformation can be found in [27]. Note that in (7) (and what follows) we use the standard notation in empirical processes theory: $\operatorname{P}(f)$ (or simply $\operatorname{P}f$) stands for the mathematical expectation of $f$ with respect to $\operatorname{P}$.
The existence of the mean embedding implies that the kernel distance in (3), as well as the supremum kernel distance in (4), are well-defined. Indeed, they are integral probability metrics; see [39]. To see this, let us consider the unit ball of $\mathcal{H}_{k}$, that is,
@@
where $(a)$ follows from the definition of mean embedding (1), $(b)$ from Pettis integrability, and $(c)$ from the reproducing property (6); see also [12, Lemma 4]. Thus, the kernel distance (3) is the integral probability metric generated by the class $\mathcal{F}_{k}$ in (9). Therefore, the supremum kernel distance (4) admits the alternative representation
@@
From the characterizations as integral probability metrics in (10) and (11), we conclude that $d_{k}$ and $d_{k,\Lambda}$ satisfy the properties of a pseudo-metric (non-negativeness, symmetry, triangular property). However, to ensure the identifiability property of a metric $d$ (i.e., $d(\operatorname{P},\operatorname{Q})=0$ if and only if $\operatorname{P}=\operatorname{Q}$) additional conditions are needed. It can be checked that when $\mathcal{X}=\mathbb{R}^{d}$, identifiability is satisfied for the usual kernels (such as the Gaussian kernel in (2)). However, when $\mathcal{X}$ is infinite-dimensional this type of results are more complicated; see [19] for a deep study of this topic for the Gaussian kernel (2). More details can also be found in [26] and [27].
@@
As a supremum over a class of functions is involved in (12), the theory of empirical processes comes into play naturally. Given a collection of functions $\mathcal{F}$, we recall that the $\mathcal{F}$-indexed empirical process (associated to $\operatorname{P}$) is $\mathbb{G}_{n}^{\operatorname{P}}=\sqrt{n}\,\left(\mathbb{P}_{n}-\operatorname{P}\right)$. The class $\mathcal{F}$ is called P-Donsker if $\mathbb{G}_{n}^{\operatorname{P}}\rightsquigarrow\mathbb{G}_{\operatorname{P}}$ in $\ell^{\infty}(\mathcal{F})$, the space of bounded real functionals defined on $\mathcal{F}$ with the supremum norm; see [48]. Here, $\rightsquigarrow$ stands for weak convergence in $\ell^{\infty}(\mathcal{F})$ and $\mathbb{G}_{\operatorname{P}}$ is a $\operatorname{P}$-Brownian bridge, that is, a zero-mean Gaussian process with covariance function